Mathematics 241 – Multivariable Calculus Quiz 3 – September 24, 2010

Your Name:

A) (10) Sketch the parametric curve α(t) = (4 cos(2t), 3 sin(2t)), for 0 ≤ t ≤ π/4. Show the starting point, the final point and indicate the direction with an arrow.

B) (10) Find a parametrization of the tangent line to the curve from part A) at the point with t = π/8.